Suppose that N= squarerootx +2 models the number of cases of an infection, in millions, of a disease x years from now.
a) Approximately how many cases of the infection will there be 16 years from now?


b) In approximately how many years will there be 7 million cases?


Answer by stanbon(75887)

Suppose that N= squareroot(x) +2 models the number of cases of an infection, in millions, of a disease x years from now.
a) Approximately how many cases of the infection will there be 16 years from now?
Let x = 16 to get:
N= sqrt(16) + 2
N= 4 + 2
N= 6 million
==================
b) In approximately how many years will there be 7 million cases?
Let N= 7 and solve for "x":
7 = sqrt(x) + 2
sqrt(x) = 5
x = 25 yrs.
=====================
Cheers,
Stan H.




N= squarerootx +2
---------
Is the equation N= sqrt(x) + 2
or is it N=sqrt(x+2) ??
Cheers,
Stan H.